OpenStudy (anonymous):

Simplify the expression 5^3 × 5^8. and explain how please

OpenStudy (anonymous):

Here is the general rule: \[ a^b*a^c=a^{b+c}\]

OpenStudy (anonymous):

\[x^a\cdot x^b=x^{a+b}\] Laws of Exponents

OpenStudy (isaiah.feynman):